OpenStudy (anonymous):

Differential Equation: x^2 y' + y = x^2 e^(-x) What I'v done: y' + 1/x^2 * y = e^(-x) Integrating factor: e^(-1/x) Which leads to: Integral of: e^(-x) * e^(-1/x) Wolfram says that there are no elementary function for this integral. Solving the homogeneous eqn: y' + 1/x^2 y = 0 Gives y_h = Ce^(1/x) which is correct as far as I know. But how can I find the particular solution?

OpenStudy (anonymous):

wolframalpha is right

OpenStudy (anonymous):

Your y_h solution is right. The particular solution is given by a non-elementary integral \[ y_p=e^{\frac{1}{x}} \int_1^x e^{-x-\frac{1}{x}} \, dx \]

OpenStudy (anonymous):

\[ y_p=e^{\frac{1}{x}} \int_1^x e^{-t-\frac{1}{t}} \, dt \]

OpenStudy (anonymous):

May be you should try Series solution of it

OpenStudy (anonymous):

Here a series solution up to power of 10 \[y(x)=x^2-3 x^3+\frac{19 x^4}{2}-\frac{229 x^5}{6}+\frac{1527 x^6}{8}-\frac{137431 x^7}{120}+\\ \frac{5772103 x^8}{720}-\frac{107745923 x^9}{1680}+\\ \frac{23273119369 x^{10}}{40320}+O\left(x^{11}\right)\]

OpenStudy (anonymous):

The above solution was generated by mathematica using the script below n = 10; y[x_] = Sum[a[i] x^i, {i, 0, n}] + O[x]^(n + 1); eqns = LogicalExpand[x^2 y'[x] + y[x] == x^2 E^-x]; ans = Solve[eqns]; g[x_] = y[x] /. ans[[1]]

OpenStudy (anonymous):

Thank you! But there exists a simple solution: Ce^(1/x) + x*e^(1/x) => y_p = x * e^(1/x) y_p satisfies the equation but I can't figure out how the get it!